OpenStudy (anonymous):

What is the sum of the n terms of the series 5 + 10 + 20 + 40 + ... ? Hint: Sn = A1(1 - r^n)/1 - r, r = 1, where A1 is the first term and r is the common ratio.

OpenStudy (anonymous):

\[Sn = 5(2^n-1)/n-1\]

OpenStudy (anonymous):

\[Sn= n/2(5+5n)\]

OpenStudy (anonymous):

\[sn=5(1-2n)\]

OpenStudy (anonymous):

\[sn=5(2^n - 1)\]

OpenStudy (anonymous):

the common ratio, r = 2. \[S_n=\frac{ 5(1-2^n)}{ 1-2 }=-5(1-2^n)\]